# Systematical study of exotic $1^{-+}$ tetraquark spectroscopy

### v1

@@ III Results and Discussion
After testing the convergence of the mass calculation for the three tetraquark configurations, light ($qq\bar{q}\bar{q}$), charmonium-like ($qc\bar{q}\bar{c}$), and fully-charmed ($cc\bar{c}\bar{c}$), we use $n=6$ as the principal quantum number for all cases. The choice yields stable results and corresponds to a basis size of $N=13$. The masses obtained with this setting are taken as the final results in this section.
@@ III.1 Results of light tetraquark
The mass results for the mixed light tetraquark states are given in table 7, where En (n = 1,2,3,…) represents the $n$th eigenstate of the Hamiltonian in Eq. (II.2) with color-spin configuration mixing. The branching ratios of the E1 to E6 with various mixed configurations are also listed in Tables 7. THR represents the threshold of two meson final states $M_{1}+M_{2}$, with the units of MeV.
Table 7: Masses for mixed $qq\bar{q}\bar{q}$ tetraquark. The mass results are given in units of MeV. THR represents the threshold of two meson final states $M_{1}+M_{2}$, with the units of MeV.
@@
$M_{cal}$ $I=0$ 1953 2210 2233 2307 2504
$\eta f_{1}(1285)$ 1835 2 <1<1 1 54 16
$\omega h_{1}(1170)$ 1952 <1<1 12 273 117
$M_{cal}$ $I=1$ 1880 1920 2104 2239 2264
$\pi f_{1}(1285)$ 1420 49 44 1 <1<1 <1<1
$\rho h_{1}(1170)$ 1940 1 1 1
For an S-wave decay of an exotic $J^{PC}=1^{-+}$ state, the only two-meson combinations allowed by the quantum numbers are $\left(0^{-+}\oplus 1^{++}\right)$ and $\left(1^{--}\oplus 1^{+-}\right)$. All other $S+P$ (S-wave + P-wave meson) combinations either fail to satisfy $C=+$ or cannot produce $J=1$ when the relative orbital angular momentum between the two mesons is $L_{rel}=0$.
It is worth mentioning that, in principle, the $I=0$ $1^{-+}$ tetraquark can decay through the $\eta\eta^{\prime}$ channel when the relative orbital angular momentum is $L_{rel}=1$, which allows $J=1$ and negative parity $P=$-. For a two-meson final state composed of two different self-conjugate mesons $M_{1}\neq M_{2}$ (each with a well-defined intrinsic $C_{M_{1}},\,C_{M_{2}}=\pm 1$), the total two-body C-parity is simply $C_{tot}=C_{M_{1}}C_{M_{2}}$, because there is no interchange symmetry between the two distinct mesons. Thus, for the $\eta\eta^{\prime}$ channel, the charge conjugation can take $C_{tot}=C_{M_{1}}C_{M_{2}}=+$ Shi et al. (2024a); Matveev et al. (2025).
@@ III.2 $\pi_{1}(2015)$
VES also performs the similar partial-wave analysis, and the $\pi_{1}(1600)$ is observed in both $f_{1}(1285)\pi$ and $b_{1}(1235)\pi$ decay channels Amelin et al. (2005). Their intensity of the $1^{-+}$ partial wave in both channel above 1.9 GeV is considered to be consistent with that observed by E852 Meyer and Swanson (2015). However, VES did not comment on this or claimed the existence of the $\pi_{1}(2015)$.
COMPASS recently presents a prelimnary results of partial-wave decomposition of the diffractive reaction $\pi^{-}p\to\pi^{-}\pi^{+}\pi^{-}\eta p$ Spülbeck (2025), and no clear evidence for an excited state as previously reported by E852 Kuhn et al. (2004) is found, and they highlight the model dependence of such interpretations.
@@
The observed S+PS{+}P decay pattern matches expectations for hybrids, yet is also natural for a compact tetraquark. a QCD sum rule analysis of exotic tetraquark states, using diquark-antidiquark currents with $I^{G}J^{PC}=1^{-}1^{-+}$ and $0^{+}1^{-+}$Chen et al. (2008a, b) can accommodate $\pi_{1}(2015)$, with the $qs\bar{q}\bar{s}$ configuration around $2.0~\text{GeV}$, which supports the assignment of $\pi_{1}(2015)$ to a compact tetraquark that naturally prefers S+PS{+}P decays such as $b_{1}\pi$ and $f_{1}\pi$.
@@ III.3 $\eta_{1}(1855)$
In 2022, the BESIII Collaboration reported the discovery of a new isoscalar exotic state, $\eta_{1}(1855)$ with quantum numbers $J^{PC}=1^{-+}$, observed in the decay process $J/\psi\to\gamma\eta\eta^{\prime}$ Ablikim et al. (2022b). The measured mass is $(1855\pm 9^{+6}_{-1})$ MeV, and the width is $(188\pm 18^{+3}_{-8})$ MeV, with a statistical significance greater than $19\sigma$.
@@
In the hybrid meson framework, the observation of $\eta_{1}(1855)$, together with the isovector $\pi_{1}(1600)$, provides valuable information for understanding the $1^{-+}$ hybrid nonet Qiu and Zhao (2022); Chen et al. (2022); Shastry et al. (2022); Wang et al. (2022); Chen et al. (2023a); Swanson (2023); Chen et al. (2023b); Shastry and Giacosa (2023). A semirelativistic quark potential model was used to estimate the masses of light hybrid mesons with $J^{PC}=1^{-+}$ Chen et al. (2023b). This study suggests that $\pi_{1}(1600)$ and $\eta_{1}(1855)$ might correspond to the hybrid configurations $(u\bar{u}-d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$, respectively.
@@
On the other hand, a study using the chiral quark model did not support the $K\bar{K}_{1}$ molecular interpretation of $\eta_{1}(1855)$ Tan et al. (2025). In another analysis using QCD sum rules Liu et al. (2025), the exotic $K\bar{K}_{1}(1270)$ and $K\bar{K}_{1}(1400)$ molecular states with $J^{PC}=1^{-+}$ were investigated, but the spectral functions turned out negative in physical regions, making the results unreliable.
@@
In the work, the theoretical results of branching ratio for $1^{-+}$ light tetraquark state decaying to $\eta\eta^{\prime}$ are all equal to zero. Thus, the $\eta_{1}(1855)$ state observed from $\eta\eta^{\prime}$ Ablikim et al. (2022b) cannot be accommodated in the current tetraquark picture, and the state could be other structures, for example, a hadronic molecule state, or a hybrid meson as reviewed above. While some models support its exotic structure, others challenge these assignments, highlighting the need for further theoretical and experimental studies.
III.4 Results of charmonium-like $1^{-+}$ tetarquark
@@
At present, no charmonium-like state with $J^{PC}=1^{-+}$ has been confirmed. Extensive studies of the $XYZ$ sector by BESIII, LHCb, Belle, and Belle II have found many states with non-exotic, non-$1^{-+}$ quantum numbers (e.g. $1^{++},1^{+-},0^{++},2^{++}$), while searches for $1^{-+}$ have so far found nothing. A dedicated BESIII search, $e^{+}e^{-}\to\gamma\eta^{(\prime)}\eta_{c}$ at $\sqrt{s}=4.258\text{-- }4.681~\mathrm{GeV}$, saw no significant signal and set 90% C.L. upper limits on the cross sections Ablikim et al. (2025).
@@
The mass spectrum and S-wave decay ratio are listed in Table 8 for the charmonium-like tetraquark. The lowest mass is around 4.2 GeV with a relative large ratio to $\eta\chi_{c1}$ channel. In the work, the theoretical results of branching ratio for $1^{-+}$ charmonium-like tetraquark state decaying to $\eta\eta_{c}$ are very small. The work suggests searching for tetraquark states in $\eta\chi_{c1}$ channels in the mass range of 4.24.34.2-4.3 GeV due to the large decay ratio of two tetraquark states.
Table 8: Masses for mixed tetraquark. The theoretical mass results $M_{cal}$ are given in units of MeV. THR represents the threshold of two meson final states $M_{1}+M_{2}$, with the units of MeV.
@@
$M_{cal}$ 4237 4279 4281 4300 4415 4435
$\rho h_{c}$ 4295 5 <<1 1
$\eta\chi_{c1}$ 4060 18 24 1 4 <<1 <<1
III.5 Results of fully-charm $1^{-+}$ tetarquark
@@
Table 9: Masses for $1^{-+}$ $cc\bar{c}\bar{c}$ tetraquark. The mass results are given in units of MeV. THR represents the threshold of two meson final states $M_{1}+M_{2}$, with the units of MeV.
@@
$M_{cal}$ 6586 6754 6870 6903 7059 7074
$J/\psi h_{c}$ 6622 1 10 310 247 1
$\eta_{c}\chi_{c1}$ 6496 54 1 10 310 247 1
The mass spectra and S-wave decay ratio are listed in Tables 9 for the fully-charmed tetraquark. The lowest mass is around 6.6 GeV with a relative large ratio to $\eta_{c}\chi_{c1}$ channel. The work suggests searching for tetraquark states in $\eta_{c}\chi_{c1}$ channels around 6.6 GeV, and in $J/\psi h_{c}$ and $\eta_{c}\chi_{c1}$ channels in the mass range of 6.9-7.1 GeV, due to the large decay ratio of several tetraquark states.
@@ IV Summary
The masses and decay branching ratio of $1^{-+}$ light, charmonium-like, and fully-charm tetraquark states are studied using a constituent quark model (CQM) that incorporates a Cornell-like potential and the Breit-Fermi interaction. The theoretical results for tetraquarks are compared with those of $1^{-+}$ exotic states.
The lowest $1^{-+}$ light tetraquark state is around 1.9 GeV, which is significantly far from the observed $\pi_{1}(1400)$ and $\pi_{1}(1600)$ states.
The observed states $\eta_{1}(1855)$ and $\pi_{1}(2015)$ are located within the light tetraquark range. The theoretical results of branching ratio for $1^{-+}$ Y tetraquark state decaying to $\eta\eta^{\prime}$ are equal to zero. Thus, the $\eta_{1}(1855)$ state observed from $\eta\eta^{\prime}$ by BESIII cannot be accommodated in the current tetraquark picture.
@@
The lowest $1^{-+}$ fully-charmed tetraquark state is around 6.6 GeV and may be searched in $\eta_{c}\chi_{c1}$ channel.

### v2

@@ III Results and Discussion
We tested convergence with increasing bases and adopt $N_{\max}=13$ (corresponding to $n_{\max}=6$ for $L=1$), which yields stable masses in all three tetraquark configurations. The masses obtained with this setting are taken as the final results in this section.
@@ III.1 Results of light tetraquark
The mass results for the mixed light tetraquark states are given in Table 7, where En (n = 1,2,3,…) represents the $n$th eigenstate of the Hamiltonian in Eq. (II.2) with color-spin configuration mixing. The relative decay ratios of the E1 to E5 with various mixed configurations are also listed in Table 7, which are normalized to the E3 $\omega h_{1}$ channel =1=1 for $I=0$ states and E3 $\rho h_{1}$ channel =1=1 for $I=1$ states. THR represents the threshold of two meson final states $M_{1}+M_{2}$, with the units of MeV. The symbol “…” in table indicates that the predicted mass lies below the corresponding two-body threshold; therefore this decay channel is kinematically forbidden in our framework.
Table 7: Masses of mixed $qq\bar{q}\bar{q}$ tetraquark states. The mass results are given in units of MeV. THR denotes the threshold of the two-meson final state $M_{1}+M_{2}$, also in units of MeV. The decay ratios are normalized to the E3 $\omega h_{1}$ channel (=1=1) for $I=0$ states and to the E3 $\rho h_{1}$ channel (=1=1) for $I=1$ states.
@@
$M_{cal}$ $I=0$ 1961 2231 2242 2320 2516
$\omega h_{1}(1170)$ 1952 0.01 0.01 1 32 26
$\eta f_{1}(1285)$ 1835 0.03 0.01 0.1 7 4
$M_{cal}$ $I=1$ 1884 1924 2112 2247 2270
$\rho h_{1}(1170)$ 1940 1 1.6 2.0
$\pi f_{1}(1285)$ 1420 38 42 2.0 1.2 0.2
For an S-wave decay of an exotic $1^{-+}$ state, the only two-meson combinations allowed by the quantum numbers are $\left(0^{-+}\oplus 1^{++}\right)$ and $\left(1^{--}\oplus 1^{+-}\right)$. All other $S+P$ (S-wave + P-wave meson) combinations either fail to satisfy $C=+$ or cannot produce $J=1$ when the relative orbital angular momentum between the two mesons is $L_{rel}=0$.
It is worth mentioning that, in principle, the $I=0$ $1^{-+}$ tetraquark can decay through the $\eta\eta^{\prime}$ channel when the relative orbital angular momentum is $L_{rel}=1$, which allows $J=1$ and negative parity $P=$-. For a two-meson final state composed of two different self-conjugate mesons $M_{1}\neq M_{2}$ (each with a well-defined intrinsic $C_{M_{1}},\,C_{M_{2}}=\pm 1$), the total two-body C-parity is simply $C_{tot}=C_{M_{1}}C_{M_{2}}$, because there is no interchange symmetry between the two distinct mesons. Since $\eta$ and $\eta^{\prime}$ are distinct, no exchange symmetry factor applies. Thus, for the $\eta\eta^{\prime}$ channel, the charge conjugation can take $C_{tot}=C_{M_{1}}C_{M_{2}}=+$ Shi et al. (2024a); Matveev et al. (2025).
@@ III.2 $\pi_{1}(2015)$
VES performed a similar partial-wave analysis, and the $\pi_{1}(1600)$ is observed in both $f_{1}(1285)\pi$ and $b_{1}(1235)\pi$ decay channels Amelin et al. (2005). Their intensity of the $1^{-+}$ partial wave in both channel above 1.9 GeV is considered to be consistent with that observed by E852 Meyer and Swanson (2015). However, VES did not comment on this and did not claim the existence of the $\pi_{1}(2015)$.
COMPASS recently presented preliminary results of partial-wave decomposition of the diffractive reaction $\pi^{-}p\to\pi^{-}\pi^{+}\pi^{-}\eta p$ Spülbeck (2025), and no clear evidence for an excited state as previously reported by E852 Kuhn et al. (2004) was found, and they highlighted the model dependence of such interpretations.
@@
[…]
The observed $S+P$ decay pattern matches expectations for hybrids, yet is also natural for a compact tetraquark. A QCD sum rule analysis of exotic tetraquark states, using diquark-antidiquark currents with $I^{G}J^{PC}=1^{-}1^{-+}$ and $0^{+}1^{-+}$ Chen et al. (2008a, b) can accommodate $\pi_{1}(2015)$, with the $qs\bar{q}\bar{s}$ configuration around $2.0~\text{GeV}$, which supports the assignment of $\pi_{1}(2015)$ to a compact tetraquark that naturally prefers $S+P$ decays such as $b_{1}\pi$ and $f_{1}\pi$.
@@ III.3 $\eta_{1}(1855)$
In 2022, the BESIII Collaboration reported the discovery of a new isoscalar exotic state, $\eta_{1}(1855)$ with quantum numbers $1^{-+}$, observed in the decay process $J/\psi\to\gamma\eta\eta^{\prime}$ Ablikim et al. (2022b). The measured mass is $(1855\pm 9^{+6}_{-1})$ MeV, and the width is $(188\pm 18^{+3}_{-8})$ MeV, with a statistical significance greater than $19\sigma$.
@@
In the hybrid meson framework, the observation of $\eta_{1}(1855)$, together with the isovector $\pi_{1}(1600)$, provides valuable information for understanding the $1^{-+}$ hybrid nonet Qiu and Zhao (2022); Chen et al. (2022); Shastry et al. (2022); Wang et al. (2022); Chen et al. (2023a); Swanson (2023); Chen et al. (2023b); Shastry and Giacosa (2023). A semirelativistic quark potential model was used to estimate the masses of light hybrid mesons with $1^{-+}$ Chen et al. (2023b). This study suggests that $\pi_{1}(1600)$ and $\eta_{1}(1855)$ might correspond to the hybrid configurations $(u\bar{u}-d\bar{d})g/\sqrt{2}$ and $s\bar{s}g$, respectively.
@@
On the other hand, a study using the chiral quark model did not support the $K\bar{K}_{1}$ molecular interpretation of $\eta_{1}(1855)$ Tan et al. (2025). In another analysis using QCD sum rules Liu et al. (2025), the exotic $K\bar{K}_{1}(1270)$ and $K\bar{K}_{1}(1400)$ molecular states with $1^{-+}$ were investigated, but the spectral functions turned out negative in physical regions, making the results unreliable.
@@
In this work, the theoretical relative decay ratios for the decay of the $1^{-+}$ light tetraquark state into $\eta\eta^{\prime}$ are found to be nearly zero. Consequently, the $\eta_{1}(1855)$ state observed in the $\eta\eta^{\prime}$ channel Ablikim et al. (2022b) cannot be accommodated within the present tetraquark picture when both the mass spectrum and decay properties are taken into account.
The $\eta_{1}(1855)$ may instead correspond to other configurations, such as a hadronic molecular state or a hybrid meson, as discussed above. In particular, if the $\eta_{1}(1855)$ is interpreted as a $K_{1}\bar{K}$ hadronic molecule and decays into $\eta\eta^{\prime}$, the corresponding decay width is expected to be sizable, as discussed in detail in Ref. Yan et al. (2023). While some models support such exotic interpretations, others challenge these assignments, highlighting the need for further theoretical and experimental studies.
III.4 Results of charmonium-like $1^{-+}$ tetraquark
@@
At present, no charmonium-like state with $1^{-+}$ has been confirmed. Extensive studies of the $XYZ$ sector by BESIII, LHCb, Belle, and Belle II have found many states with non-exotic, non-$1^{-+}$ quantum numbers (e.g. $1^{++},1^{+-},0^{++},2^{++}$), while searches for $1^{-+}$ have so far found nothing. A dedicated BESIII search, $e^{+}e^{-}\to\gamma\eta^{(\prime)}\eta_{c}$ at $\sqrt{s}=4.258\text{-- }4.681~\mathrm{GeV}$, saw no significant signal and set 90% C.L. upper limits on the cross sections Ablikim et al. (2025).
@@
The mass spectrum and S-wave relative decay ratios of the charmonium-like tetraquarks are listed in Table 8, and the ratios are normalized to the E3 $\eta\chi_{c1}$ channel =1=1. Unlike the cases for light tetraquarks and fully-charm tetraquarks, where the flavor configurations impose constraints on color-spin-orbit ones, the color-spin-orbit configurations for charmonium-like tetraquark are independent of flavors. Therefore, the present work is isospin-blind for charmonium-like tetraquark states. As a result, the masses of the $I=0$ and $I=1$ states are the same in this work, and both the $\pi/\eta+\chi_{c1}$ and $\rho/\omega+h_{c}$ decay channels are included. The relative decay ratios differ due to phase-space factors.
The lowest mass is around 4.2 GeV with a relative large ratio to $\eta\chi_{c1}$ channel. In the work, the theoretical results of relative decay ratios for $1^{-+}$ charmonium-like tetraquark state decaying to $\eta\eta_{c}$ are very small. The work suggests searching for tetraquark states in $\eta\chi_{c1}$ channels in the mass range of 4.24.34.2-4.3 GeV due to the large decay ratio of two tetraquark states.
Table 8: Masses of mixed charmonium-like tetraquark states. The theoretical mass results $M_{cal}$ are given in units of MeV. THR denotes the threshold of the two-meson final state $M_{1}+M_{2}$, also in units of MeV. The decay ratios are normalized to the E3 $\eta\chi_{c1}$ channel (=1=1).
@@
$M_{cal}$ $I=0$ 4245 4287 4289 4307 4421 4441
$\eta\chi_{c1}$ 4057 13 23 1 4.5 0.1 0.6
$\omega h_{c}$ 4307 0.03 0.3 0.6
$M_{cal}$ $I=1$ 4245 4287 4289 4307 4421 4441
$\pi\chi_{c1}$ 3650 15 21 0.01 0.01 0.1 0.8
$\rho h_{c}$ 4295 0.2 0.3 0.7
III.5 Results of fully-charm $1^{-+}$ tetraquark
@@
Table 9: Masses of $1^{-+}$ $cc\bar{c}\bar{c}$ tetraquark states. The mass results are given in units of MeV. THR denotes the threshold of the two-meson final state $M_{1}+M_{2}$, also in units of MeV. The decay ratios are normalized to the E3 $J/\psi h_{c}$ channel (=1=1).
@@
$M_{cal}$ 6595 6762 6881 6908 7063 7079
$J/\psi h_{c}$ 6622 0.1 1 36 47 0.2
$\eta_{c}\chi_{c1}$ 6496 2 0.1 1.5 51 60 0.3
The mass spectra and S-wave relative decay ratio are listed in Tables 9 for the fully-charm tetraquark, and the ratios are normalized to the E3 $J/\psi h_{c}$ channel =1=1. The lowest mass is around 6.6 GeV with a relative large ratio to $\eta_{c}\chi_{c1}$ channel. The work suggests searching for tetraquark states in $\eta_{c}\chi_{c1}$ channels around 6.6 GeV, and in $J/\psi h_{c}$ and $\eta_{c}\chi_{c1}$ channels in the mass range of 6.9-7.1 GeV, due to the large decay ratio of several tetraquark states.
@@ IV Summary
The masses and decay patterns of $1^{-+}$ light, charmonium-like, and fully-charm tetraquark states are studied using a constituent quark model that incorporates a Cornell-like potential and the Breit-Fermi interaction. The theoretical results for tetraquarks are compared with those of $1^{-+}$ exotic states.
The lowest $1^{-+}$ light tetraquark state is around 1.9 GeV, which is significantly above the observed $\pi_{1}(1400)$ and $\pi_{1}(1600)$ states.
The observed states $\eta_{1}(1855)$ and $\pi_{1}(2015)$ are located within the light tetraquark range. $1^{-+}$ Y tetraquark states decaying to $\eta\eta^{\prime}$ is strongly suppressed in our CSF overlap within the rearrangement mechanism. Thus, the $\eta_{1}(1855)$ state observed from $\eta\eta^{\prime}$ by BESIII cannot be accommodated in the current tetraquark picture.
@@
The lowest $1^{-+}$ fully-charm tetraquark state is around 6.6 GeV and may be searched in $\eta_{c}\chi_{c1}$ channel.